User Tools

Site Tools


operations:agenda2018_04_19

Chair: Katie

Attended: Jonny, Ross, Atifur, Tiege, Mas, Simin, Lucas, Patrick, Lucia, Jamie

Apologies: Arwin

Absent:

Handover Issues

  • R4836:
    • yg: e-remote control lost connection. In the end Jamie fixed the problem, as Yg weren't answering the phone after the module change.
    • 1954 extension doesnt work if connection isnt up, the other phone should work.
    • ADSL router had started blocking vpn packets. Brett fixed it by turning off adsl connection…he is still looking at it. THere is a new router which will hopefully help.
  • APSG42:
    • Hobart: the Field system time is 0.5 seconds off mk5 time.
    • field system time not running on indp time goes out of allignment by a fraction of a second. Means we start and stop recording a little off the sched. Possible fix see below.
    • Hobart: No autocorrelations.“ I think because I manually drudged the proc file for Hobart26m. Could “hack” the proc file with that from the previous apsg containing Hobart26m”. Copied a previous checkmk5 from a aua039 proc file. checkmk5. (worked?) - should add fix to wiki
    • Hobart: Last BBC in S band looks pretty dodgy.
    • Hobart: The 70K stage temperature fluctuated a lot, every few scans. Is this normal? Sensor that its picking up is becoming unstable, so wrong but not dramatically wrong.
    • Hobart: formatter to fs time difference 0.5 seconds or greater' error, fixed with 'sy=run setcl offset' as suggested in wiki, while not recording (doesnt introduce a clock jump).
  • R4837:
    • Ke: Network connection lost to pcfs, mk5, time etc
    • Ke: Mark 5 crashed bringing down network. Connection restored on-site. The whole network had died, probably a power glitch there. Its being replaced with a (hopefully) good one.
  • R1838:
    • Ke:Large different between formatter and maserdelay setting off Alarm. Delay through DBBC has changed, monitor for stability and reset DBBC if it drifts. Had to reconfigure and restart the DBBC several times.c Mas called yg they said that they fixed it, and they had access. Here the vpn connection wasnt coming up. Brett had to switch off the router and wait for connection to come back.
    • Yg: Network connection lost (again). Restarted DBBC.
    • Yg: VNC failed. Brett had to reset.
  • AUA042:
    • Hobart: Live page frozen - restart the telmon web server.
    • Ke: Delay Drifting. Need to reconfigure/restart the DBBC. Didnt work, so rebooted whole computer and that fixed it.
    • Slewing errors - again with the schedule file - not as bad as last time. Just beeps, no alarms.

General Issues

  
  *for AUA042 the disk pos entry isn't working. have to manually figure them out.
  *Stowing Yaragadee should be fixed at the moment. Stowing should be fine.
 
/home/www/auscope/opswiki/data/pages/operations/agenda2018_04_19.txt · Last modified: 2018/04/19 03:24 by Katie Vandorou